OpenStudy (science0229):

Can someone help me with a combinatorics problem? I want to give a combinatorial proof, but I'm stuck from the beginning!

OpenStudy (science0229):

The problem is to prove the following. \[\sum_{k=0}^{n}\left(\begin{matrix}n+k \\ k\end{matrix}\right)\left( \frac{ 1 }{ 2 } \right)^k=2^n\]

OpenStudy (science0229):

@welshfella

OpenStudy (science0229):

I proved it using mathematical induction, but can't think of any way to give a combinatorial proof("proof by counting")...

OpenStudy (science0229):

@mathstudent55